# Non-monotonic specific entropy on the transition line near the QCD critical point

### v1

@@ III.4 Numerical values from BEST EOS
Set $\mu_{\rm c}$ (MeV) $T_{\rm c}$ (MeV)  $w$  $\rho$ $\alpha_{1}$ (${}^{\circ}$) $\alpha_{2}$ (${}^{\circ}$)
@@ 2 $-5$
Set $\Delta\mu_{1}$ (MeV) $\Delta T_{1}$ (MeV) $|r_{\text{max}}|$ $\hat{s}_{\text{max}}-\hat{s}_{\rm c}$
@@ III.5 Justification for neglecting the regular terms
Figure 4: Critical specific entropy $\hat{s}_{\rm c}$ in the parametric family of EOS from Ref. [3] as a function of the critical parameter $\mu_{\rm c}$ while $w$, $\rho$, and $\alpha_{2}$ parameters are fixed as in Table 1 for Set 1 (solid line) and Set 2 (dashed line). Black circles represent values in Eq. (36). The dotted line represents $\cot\alpha_{1}$ as a function of $\mu_{\rm c}$. Since $\hat{s}_{\rm c}>\cot\alpha_{1}$ these values support the scenario where the non-monotonic side appears on the HRG side, according to Eq. (33).
@@ IV Isentropic trajectories near the QCD critical point
In addition, this analysis uses the expression of $\phi(h,r)$ and $\varepsilon(h,r)$ determined by the critical EOS. The derivation of these can be found in Appendix B. To evaluate $\hat{s}$ on the $(\mu,T)$ plane, we first solve Eq. (II.1), leading to
@@
In the illustrations below, we use the same Sets 1 and 2 defined in Table 1. As we observed in the previous section, the values of $\hat{s}_{\rm c}$ for both Set 1 and 2 given by Eq. (36) and $\cot\alpha_{1}=14.9$ from Ref. [3] correspond to the scenario with the specific entropy being non-monotonic on the low-$T$ (HRG) side of the coexistence line, according to the criterion in Eq. (33). To illustrate and compare with the scenario where $\hat{s}$ is non-monotonic on the QGP side, we consider another alternative value, $\hat{s}_{\rm c}=12.<\cot\alpha_{1}$, for both Sets 1 and 2. Thus, for each Set 1 and 2 we now have two alternatives distinguished by the value of $\hat{s}_{\rm c}$:
$\hat{s}_{\rm c}=20.3\,,19.3\,\quad\mbox{or}\quad 12.$ (40)
@@
Figure 5: Contours of $\hat{s}$ computed using Eq. (7): (a) and (b) on $(r,\,h)$ plane; (c) and (d) on $(\mu,\,T)$ plane. We set $\hat{s}_{\rm c}$ in the first two of Eq. (40) and use two sets of parameters (see Table 1) different by the value of $\alpha_{2}$ shown at the top of each panel. Set 1 is used in panels (a) and (c), while Set 2 – in (b) and (d). The right legend summarizes the notations of points and curves used in the plots: the black, green, and gray points depict the critical point, the maximum, and the critical double along the nonmonotonic branch as shown in Fig. 3; the red and blue lines and the black dashed lines denote $\hat{r}$ axis on $h=+0$ and $h=-0$, and the $\hat{h}$ axis ($r=0$); the black and green solid curves denote the contours with $\hat{s}=\hat{s}_{\rm c}$ and $\hat{s}_{\rm max}$ given by Eq. (16); the magenta and cyan dashed curves show the ridgeline and the valley lines generated based on the mathematical definitions in Eq. (104). The ridge line intersects the coexistence line away from the critical point, even though, due to the smallness of $r_{\rm max}$, it is hard to see, especially for large $\alpha_{2}$ case [(a) and (c)], as discussed in the main text. The non-monotonic side appears on the HRG side, according to Eq. (33), for the present parameter choice.
@@
Figure 6: Same as Fig. 5 with the only difference being the choice of $\hat{s}_{c}=12$ from Eq. (40). Since ${\rm sgn}(\hat{s}_{\rm c}-\cot\alpha_{1})=-1$, the ridge and the non-monotonic behavior of $\hat{s}$ along the coexistence line is now on the QGP side, according to Eq. (33).
@@
We present the contour plots for $\hat{s}_{\rm c}=20.3\,,19.3$ and $\hat{s}_{\rm c}=12.$ in Figs. 5 and 6, respectively. For each set of the parameters shown at the top, we have depicted the contours on two distinct planes, $(r,\,h)$ and $(\mu,\,T)$, represented as panels (a), (b) and (c), (d), respectively. All the contours exhibit a distinctive ridge structure. We have drawn characteristic ridge lines and valley lines in magenta and cyan dashed lines, respectively, whose mathematical definitions are given by (104) (see Appendix D for details).
@@
The appearance of the non-monotonic structure along the coexistence line can be seen as a consequence of the coexistence line cutting across the ridge of $\hat{s}$. The mixing between $\phi$ and $\varepsilon$ in Eq. (7) for $\hat{s}$ modifies the ridgeline of $\varepsilon(h,\,r)$ from its symmetric location along the crossover line $h=0$ shown in Fig. 9 (d). The ridge line for $\hat{s}$ is bent and shifted along the coexistence line, intersecting it away from the critical point, at $r=r_{\rm max}$. Since $|r_{\rm max}|$ is small (see Table 2), the non-monotonic structure (the maximum and the critical double) is squeezed toward the critical point, especially for Set 1 with large $\alpha_{2}$ (see Sec. III.4 for explanation). It is evident that the non-monotonic shape observed along the coexistence line in Fig. 3 is a cross-section of the ridge.
@@ V.1 The slope formula
and Eqs. (II.2). Here, we used $\beta-1+\gamma=1-\alpha-\beta$ and introduced
@@
Figure 7 illustrates the relative angle computed by (V.1) on each side of the coexistence line for Set 2. The angle $\alpha_{\hat{s}}-\alpha_{1}$ is non-monotonic on the branch where $\hat{s}$ is non-monotonic. At the maximum of $\hat{s}$, at $x=x_{\rm max}$, the relative angle vanishes. In other words, $\hat{s}=\hat{s}_{\rm max}$ contour is tangential to the coexistence line. For $x>x_{\rm max}$ the relative angle increases, crossing ${}^{\circ}$ at the point where the denominator in Eq. (V.1) vanishes. At this point, the isentrope is perpendicular to the coexistence line.
Figure 7: The angle between an isentropic curve and the coexistence line at their intersection as a function of the distance of the intersection point from the critical point. The angle is found from Eq. (V.1) and plotted with $h\to+0$ branch in red and $h\to-0$ in blue for the choice of parameters corresponding to Set 2 in Table 1 (small $\alpha_{12}$) and $x_{\rm max}>0$ in Eq. (45). The inset shows the non-monotonic structure near the critical point.
@@
From the examples in Figs. 5 (a, c) and 6 (a, c) (Set 1 parameters), we see that $|r_{\rm max}|$ is so small that the resolution of the phase diagram is not sufficient to distinguish points $r=0$, $r_{\rm max}$, and $r_{\rm cd}$, characterizing the non-monotonic structure. There appear to be two isentropes (solid black) meeting at the critical point and forming a cusp. Such a cusp has indeed been observed in Refs. [3, 8] and subsequent literature, where Set 1 was used, as well as earlier, in Ref. [7], where $\alpha_{12}=90^{\circ}$ choice was made.
@@
In the next subsection, we shall zoom into the region $0>r>r_{\rm cd}$ to study how the non-monotonic structure there affects the isentropes in more detail.
@@ V.3 Classification of trajectories
We represent different types of isentropic trajectories near the coexistence line schematically in Fig. 8. Panels (a) and (b) correspond to the scenarios where the non-monotonicity appears on the HRG and QGP side, ${\rm sgn}(\hat{s}_{\rm c}-\cot\alpha_{1})=+1$ and $-1$, respectively. These sample trajectories intersect the coexistence line at a relative angle $\alpha_{\hat{s}}-\alpha_{1}$ in agreement with Eq. (V.1) (see also Fig. 7 for the non-monotonic HRG side scenario)
Figure 8: Classification of isentropic trajectories near the QCD critical point. In panel (a) the maximum of $\hat{s}$ is on the HRG side, while in panel (b) – on the QGP side. The arrows point in the direction the entropy density decreases (the system expands). The gray curve is the critical isentrope ($\hat{s}=\hat{s}_{\rm c}$). The contour $\hat{s}=\hat{s}_{\rm max}$ given by Eq. (16) is the green curve. The pink, purple, and blue curves are representative contours in each range of $\hat{s}$ shown in the corresponding color in panels (c) and (d). The dashed lines in (a) and (b) express details depending on the choice of $\alpha_{2}$, which can be seen as a difference between, e.g., Fig. 5 (c) and (d), away from the coexistence line. Panel (d) summarizes the classification based on the value of $\hat{s}$ (see text).
@@
(I) Trajectories with $\hat{s}_{\rm ini}>\hat{s}_{\rm max}$ do not encounter the first-order transition and pass through the crossover region. These are shown as pink curves in panels (a) and (b), corresponding to horizontal lines in the pink area in panel (c);
(II) Trajectories with $\hat{s}_{\rm c}<\hat{s}_{\rm ini}<\hat{s}_{\rm max}$ enter the coexistence region from one side (the non-monotonic one) and exit on the same side. These are shown as purple curves in panels (a) and (b), and correspond to horizontal lines in the purple region in (c);
(III) Trajectories with $\hat{s}_{\rm ini}<\hat{s}_{\rm c}$ enter the coexistence region from one side and exit on the opposite side [the blue curve in (a) and (b), and the blue region in (c)].
The same three classes of trajectories exist in both scenarios distinguished by the location of the specific entropy maximum – HRG or QGP side, as illustrated in Fig. 8 panels (a) and (b), respectively. The main difference between the two scenarios is the direction of the $T$ (and $\mu$) change upon crossing the coexistence line. Under the “HRG-side scenario” the temperature decreases, while under the “QGP-side scenario” the temperature increases. Both scenarios have been discussed in the literature (see, e.g., Refs. [4, 5, 14]).
@@ VI Conclusion
The main observation of this paper is that the specific entropy as a function of the distance from the critical point along the coexistence line must be non-monotonic, and specifically must exhibit a maximum. This is a robust and model-independent consequence of the following two facts. First, the universal critical behavior dictates that the specific entropy must rise on one of the two sides of the coexistence line as a function of $|T-T_{c}|$. Second, the third law of thermodynamics dictates that the specific entropy must fall to zero at $T=0$.
We find that the maximum could occur on either side of the coexistence line: high-$T$ (QGP) or low-$T$ (HRG) side. We find a criterion that determines which side is non-monotonic. It is determined by the sign of $\hat{s}_{\rm c}-\cot\alpha_{1}$, i.e., the difference between the specific entropy and the inverse slope of the coexistence line, both quantities evaluated at the critical point.
When this discriminant quantity is sufficiently small, the maximum occurs close to the critical point, where the equation of state is determined universally, up to a small number of non-universal parameters standardized by Ref. [3]. We use this regime to demonstrate analytically how the maximum moves as a function of the critical point parameters.
@@
Turning to the behavior of the isentropic trajectories in the ($\mu$, $T$) plane, we demonstrate that the maximum on the coexistence line can be viewed as a cross-section of a “ridge” of $s/n$ “landscape,” where the trajectories are contours or equal elevation lines. The ridge topography also helps explain the critical “focusing” or ”lensing” effects observed in Refs. [7, 8].
@@ Appendix A Susceptibilities in a scaling form
For our discussion along the coexistence line in Secs. III and V, we need the values of the scaling functions (A) at $h=\pm 0$, or, equivalently, $z=\pm 0$:121212 These coefficients are known as critical amplitudes and in the common notation (see, e.g., Ref. [15]): $\tilde{\varepsilon}_{r}(0)=A^{-}$, $|\tilde{\varepsilon}_{h}(0)|=\beta B$, and $\tilde{\phi}_{h}(0)=C^{-}$ (where “--” refers to $r<0$).
@@
Among four values of (A), only two are unknown. This is because the normalization of $\tilde{\phi}(\pm 0)=\pm 1$ have fixed one of them and the Maxwell relation,
@@
where the coefficients $a\approx-0.762$ and $b\approx 0.008$ are obtained through renormalization group analysis using a non-perturbative resummation [17, 18]. To distinguish the scaling functions of $z$ introduced in Eq. (11), we use the bar notation for the scaling form of $\theta$. The normalization constants $\phi_{0}$ and $h_{0}$ are fixed by the conditions $1=\phi(h=+0,r=-1)=\phi(h=1,r=0)$, which reduces to
@@
We numerically solve Eq. (B) over a specific closed range of $Y$, yielding $\theta(h,\,r)$ and $R(h,\,r)$ denoted by $Z(Y)$. Then, we substitute $Z(Y)$ into (53) to obtain $\phi(Y)$. In Fig. 9 (a)–(c), we present the contours of $R$, $\theta$, and $\phi$, respectively. As observed from (a) and (b), $R$ ($R\geq 0$) represents the distance from the critical point at the origin, while $\theta$ ($-\theta_{\rm max}\leq\theta\leq\theta_{\rm max})$ measures the distance along the constant $R$ contour from the edges of the first-order boundary ($h=0,\ r<0$). Here, $\theta=\pm\theta_{\rm max}$ corresponds to $h=\pm 0$. With this parameterization, we can describe the criticality as a function of the $R$ variable, and the thermodynamic quantities will exhibit smooth behavior in the $\theta$ variable.
@@
We will determine the scaling function $\bar{p}(\theta)$ below to match with the relation between $\phi$ and $h$, Eq. (B). With Eq. (56) as a starting point, we can calculate the equations for the magnetization $\phi$ and the Ising energy $\varepsilon$ as (5):
@@
To derive the right-hand sides of Eq. (B), we use the derivative formula for an arbitrary function, $f\equiv R^{a_{f}}\bar{f}(\theta)$, with the scaling dimension $a_{f}$:
@@
which determines the form of $\bar{p}(\theta)$. Given that this is a first-order nonhomogeneous equation, it includes both special and general solutions. The special solution is relevant to our current analysis concerning the critical exponents beyond the mean-field level: $\bar{p}(\theta)=\sum_{n=0}^{3}C_{n}(1-\theta^{2})^{n}$, where $C_{n}$ are functions of the critical exponents and the parameters in Eq. (B) (for those explicit forms, refer to Ref. [3]). The constant of integration associated with the general solution $\propto(\theta^{2}-1)^{2-\alpha}$ is determined to be zero so that any singularity does not arise in the scaling part of the pressure, $\bar{p}(\theta)$.
@@
We highlight the qualitative distinction in behaviors for $\phi$ and $\varepsilon$, in Fig. 9 (c) and (d), respectively. The former exhibits a monotonic increase as $\theta$ rises while keeping $R$ constant [also refer to Fig. 9 (a) and (b) for tracking $R$ and $\theta$]. In contrast, the convex nature of $\varepsilon(\theta)$ results in a non-monotonic ridge structure of $\varepsilon$. The distinctive ridgeline and the valley lines are illustrated in the magenta and cyan dashed lines according to the mathematical definition given by Eq. (104). Specifically, the ridgelines form a straight line with $\theta=0$ parametrized by an arbitrary $R>0$, due to symemtry of $\varepsilon$ under $h\rightarrow-h$ [^14] This is in contrast to $\hat{s}$ as we discuss at the end of Sec. IV.
@@
One can check that the the Maxwell relation (50) holds for the expressions (68b) and (68c), using Eq. (B).
@@
    We derived analytic expressions for $\phi(\theta,R)$, Eqs. (53) and (54a), and $\varepsilon(\theta,R)$, (B), respectively.
@@
    We numerically solved Eq. (B) with (54b), within a specified range of $h$ and $r$, and obtain $\theta(h,r)$ and $R(h,r)$.
@@
Here, the solution of the inverse function $\bar{h}_{2}^{-1}(z)$ restricted to (73) depends on the sign of $r$. If $r\leq 0$, $\bar{h}_{2,-}^{-1}$ chose the solution of $\bar{h}_{2}^{-1}(z)$ satisfies $1\leq|\theta|<\theta_{\rm max}$; if $r>0$, $\bar{h}_{2,+}^{-1}$ chose $10<|\theta|<$ 1. Using (74) on the $r\leq 0$ branch, we derive
@@
We can apply the same method to the Ising energy density $\varepsilon$. Eliminating $R$ from the scaling form in terms of $(\theta,R)$ variable, as given by (B), we have

### v2

@@ III.4 Numerical values from BEST EOS
Set $\mu_{\rm c}$ (MeV) $T_{\rm c}$ (MeV)  $w$  $\rho$ $\alpha_{1}$ (${\rm deg}$) $\alpha_{2}$ (${\rm deg}$)
@@ 2 $-5$
Set $\Delta\mu_{1}$ (MeV) $\Delta T_{1}$ (MeV) $|r_{\rm max}|$ $\hat{s}_{\rm max}-\hat{s}_{\rm c}$
@@ III.5 Justification for neglecting the regular terms
Figure 4: Critical specific entropy $\hat{s}_{\rm c}$ in the parametric family of EOS from Ref. [3] as a function of the critical parameter $\mu_{\rm c}$ while $w$, $\rho$, and $\alpha_{2}$ parameters are fixed as in Table 1 for Set 1 (solid line) and Set 2 (dashed line). Black circles represent values in Eq. (36). The dotted line represents $\cot\alpha_{1}$ as a function of $\mu_{\rm c}$. Since $\hat{s}_{\rm c}>\cot\alpha_{1}$ these values support the scenario where the nonmonotonic side appears on the HRG side, according to Eq. (33).
@@ IV Isentropic trajectories near the QCD critical point
In addition, this analysis uses the expression of $\phi(h,r)$ and $\varepsilon(h,r)$ determined by the critical EOS. The derivation of these can be found in Appendix B. To evaluate $\hat{s}$ on the $(\mu,T)$ plane, we first solve Eq. (2), leading to
@@
In the illustrations below, we use the same Sets 1 and 2 defined in Table 1. As we observed in the previous section, the values of $\hat{s}_{\rm c}$ for both Set 1 and 2 given by Eq. (36) and $\cot\alpha_{1}=14.9$ from Ref. [3] correspond to the scenario with the specific entropy being nonmonotonic on the low-$T$ (HRG) side of the coexistence line, according to the criterion in Eq. (33). To illustrate and compare with the scenario where $\hat{s}$ is nonmonotonic on the QGP side, we consider another alternative value, $\hat{s}_{\rm c}=12.<\cot\alpha_{1}$, for both Sets 1 and 2. Thus, for each Set 1 and 2 we now have two alternatives distinguished by the value of $\hat{s}_{\rm c}$:
$\hat{s}_{\rm c}=20.3\,,19.3\,\quad\mbox{or}\quad 12.,$ (40)
@@
Figure 5: Contours of $\hat{s}$ computed using Eq. (7): [(a) and (b)] On $(r,\,h)$ plane; [(c) and (d)] on the $(\mu,\,T)$ plane. We set $\hat{s}_{\rm c}$ in the first two of Eq. (40) and use two sets of parameters (see Table 1) different by the value of $\alpha_{2}$ shown at the top of each panel. Set 1 is used in panels (a) and (c), while Set 2 – in (b) and (d). The right legend summarizes the notations of points and curves used in the plots: the black, green, and gray points depict the critical point, the maximum, and the critical double along the nonmonotonic branch as shown in Fig. 3; the red and blue lines and the black dashed lines denote $\hat{r}$ axis on $h=+0$ and $h=-0$, and the $\hat{h}$ axis ($r=0$); the black and green solid curves denote the contours with $\hat{s}=\hat{s}_{\rm c}$ and $\hat{s}_{\rm max}$ given by Eq. (16); the magenta and cyan dashed curves show the ridgeline and the valley lines generated based on the mathematical definitions in Eq. (104). The ridge line intersects the coexistence line away from the critical point, even though, due to the smallness of $r_{\rm max}$, it is hard to see, especially for large $\alpha_{2}$ case [(a) and (c)], as discussed in the main text. The nonmonotonic side appears on the HRG side, according to Eq. (33), for the present parameter choice.
@@
Figure 6: Same as Fig. 5 with the only difference being the choice of $\hat{s}_{c}=12$ from Eq. (40). Since ${\rm sgn}(\hat{s}_{\rm c}-\cot\alpha_{1})=-1$, the ridge and the nonmonotonic behavior of $\hat{s}$ along the coexistence line is now on the QGP side, according to Eq. (33).
@@
We present the contour plots for $\hat{s}_{\rm c}=20.3\,,19.3$ and $\hat{s}_{\rm c}=12.$ in Figs. 5 and 6, respectively. For each set of the parameters shown at the top, we have depicted the contours on two distinct planes, $(r,\,h)$ and $(\mu,\,T)$, represented as panels (a) and (b) as well as (c) and (d), respectively. All the contours exhibit a distinctive ridge structure. We have drawn characteristic ridge lines and valley lines in magenta and cyan dashed lines, respectively, whose mathematical definitions are given by (104) (see Appendix D for details).
@@
The appearance of the nonmonotonic structure along the coexistence line can be seen as a consequence of the coexistence line cutting across the ridge of $\hat{s}$. The mixing between $\phi$ and $\varepsilon$ in Eq. (7) for $\hat{s}$ modifies the ridgeline of $\varepsilon(h,\,r)$ from its symmetric location along the crossover line $h=0$ shown in Fig. 9 (d). The ridge line for $\hat{s}$ is bent and shifted along the coexistence line, intersecting it away from the critical point, at $r=r_{\rm max}$. Since $|r_{\rm max}|$ is small (see Table 2), the nonmonotonic structure (the maximum and the critical double) is squeezed toward the critical point, especially for Set 1 with large $\alpha_{2}$ (see Sec. III.4 for explanation). It is evident that the nonmonotonic shape observed along the coexistence line in Fig. 3 is a cross section of the ridge.
@@ V.1 The slope formula
and Eqs. (6). Here, we used $\beta-1+\gamma=1-\alpha-\beta$ and introduced
@@
Figure 7 illustrates the relative angle computed by (V.1) on each side of the coexistence line for Set 2. The angle $\alpha_{\hat{s}}-\alpha_{1}$ is nonmonotonic on the branch where $\hat{s}$ is nonmonotonic. At the maximum of $\hat{s}$, at $x=x_{\rm max}$, the relative angle vanishes. In other words, $\hat{s}=\hat{s}_{\rm max}$ contour is tangential to the coexistence line. For $x>x_{\rm max}$ the relative angle increases, crossing 90 at the point where the denominator in Eq. (V.1) vanishes. At this point, the isentrope is perpendicular to the coexistence line.
Figure 7: The angle between an isentropic curve and the coexistence line at their intersection as a function of the distance of the intersection point from the critical point. The angle is found from Eq. (V.1) and plotted with $h\to+0$ branch in red and $h\to-0$ in blue for the choice of parameters corresponding to Set 2 in Table 1 (small $\alpha_{12}$) and $x_{\rm max}>0$ in Eq. (45). The inset shows the nonmonotonic structure near the critical point.
@@
From the examples in Figs. 5(a) and 5(c) and Figs. 6(a) and 6(c) (Set 1 parameters), we see that $|r_{\rm max}|$ is so small that the resolution of the phase diagram is not sufficient to distinguish points $r=0$, $r_{\rm max}$, and $r_{\rm cd}$, characterizing the nonmonotonic structure. There appear to be two isentropes (solid black) meeting at the critical point and forming a cusp. Such a cusp has indeed been observed in Refs. [3, 8] and subsequent literature, where Set 1 was used, as well as earlier, in Ref. [7], where $\alpha_{12}=90^{\circ}$ choice was made.
@@
In the next subsection, we shall zoom into the region $0>r>r_{\rm cd}$ to study how the nonmonotonic structure there affects the isentropes in more detail.
@@ V.3 Classification of trajectories
We represent different types of isentropic trajectories near the coexistence line schematically in Fig. 8. Figures 8(a) and 8(b) correspond to the scenarios where the nonmonotonicity appears on the HRG and QGP side, ${\rm sgn}(\hat{s}_{\rm c}-\cot\alpha_{1})=+1$ and $-1$, respectively. These sample trajectories intersect the coexistence line at a relative angle $\alpha_{\hat{s}}-\alpha_{1}$ in agreement with Eq. (V.1) (see also Fig. 7 for the nonmonotonic HRG side scenario)
Figure 8: Classification of isentropic trajectories near the QCD critical point. In panel (a) the maximum of $\hat{s}$ is on the HRG side, while in panel (b) it is on the QGP side. The arrows point in the direction the entropy density decreases (the system expands). The gray curve is the critical isentrope ($\hat{s}=\hat{s}_{\rm c}$). The contour $\hat{s}=\hat{s}_{\rm max}$ given by Eq. (16) is the green curve. The pink, purple, and blue curves are representative contours in each range of $\hat{s}$ shown in the corresponding color in panels (c) and (d). The dashed lines in (a) and (b) express details depending on the choice of $\alpha_{2}$, which can be seen as a difference between, e.g., Figs. 5(c) and 5(d), away from the coexistence line. Panel (d) summarizes the classification based on the value of $\hat{s}$ (see text).
@@
(I) Trajectories with $\hat{s}_{\rm ini}>\hat{s}_{\rm max}$ do not encounter the first-order transition and pass through the crossover region. These are shown as pink curves in Figs. 8(a) and 8(b), corresponding to horizontal lines in the pink area in Fig. 8(c);
(II) Trajectories with $\hat{s}_{\rm c}<\hat{s}_{\rm ini}<\hat{s}_{\rm max}$ enter the coexistence region from one side (the nonmonotonic one) and exit on the same side. These are shown as purple curves in Figs. 8(a) and 8(b), and correspond to horizontal lines in the purple region in Fig. 8(c);
(III) Trajectories with $\hat{s}_{\rm ini}<\hat{s}_{\rm c}$ enter the coexistence region from one side and exit on the opposite side [the blue curve in Figs. 8(a) and 8(b), and the blue region in Fig. 8(c)].
The same three classes of trajectories exist in both scenarios distinguished by the location of the specific entropy maximum – HRG or QGP side, as illustrated in Figs. 8(a) and 8(b), respectively. The main difference between the two scenarios is the direction of the $T$ (and $\mu$) change on crossing the coexistence line. Under the “HRG-side scenario” the temperature decreases, while under the “QGP-side scenario” the temperature increases. Both scenarios have been discussed in the literature (see, e.g., Refs. [4, 5, 14]).
@@ VI Conclusion
The main observation of this paper is that the specific entropy as a function of the distance from the critical point along the coexistence line must be nonmonotonic, and specifically must exhibit a maximum. This is a robust and model-independent consequence of the following two facts. First, the universal critical behavior dictates that the specific entropy must rise on one of the two sides of the coexistence line as a function of $|T-T_{c}|$. Second, the third law of thermodynamics dictates that the specific entropy must fall to zero at $T=0$.
We find that the maximum could occur on either side of the coexistence line: high-$T$ (QGP) or low-$T$ (HRG) side. We find a criterion that determines which side is nonmonotonic. It is determined by the sign of $\hat{s}_{\rm c}-\cot\alpha_{1}$, i.e., the difference between the specific entropy and the inverse slope of the coexistence line, both quantities evaluated at the critical point.
When this discriminant quantity is sufficiently small, the maximum occurs close to the critical point, where the equation of state is determined universally, up to a small number of nonuniversal parameters standardized by Ref. [3]. We use this regime to demonstrate analytically how the maximum moves as a function of the critical point parameters.
@@
Turning to the behavior of the isentropic trajectories in the ($\mu$, $T$) plane, we demonstrate that the maximum on the coexistence line can be viewed as a cross section of a “ridge” of $s/n$ “landscape,” where the trajectories are contours or equal elevation lines. The ridge topography also helps explain the critical “focusing” or ”lensing” effects observed in Refs. [7, 8].
@@ Appendix A Susceptibilities in a scaling form
For our discussion along the coexistence line in Secs. III and V, we need the values of the scaling functions (48) at $h=\pm 0$, or, equivalently, $z=\pm 0$:121212 These coefficients are known as critical amplitudes and in the common notation (see, e.g., Ref. [15]): $\tilde{\varepsilon}_{r}(0)=A^{-}$, $|\tilde{\varepsilon}_{h}(0)|=\beta B$, and $\tilde{\phi}_{h}(0)=C^{-}$ (where “--” refers to $r<0$).
@@
Among four values of (49), only two are unknown. This is because the normalization of $\tilde{\phi}(\pm 0)=\pm 1$ have fixed one of them and the Maxwell relation,
@@
where the coefficients $a\approx-0.762$ and $b\approx 0.008$ are obtained through renormalization group analysis using a nonperturbative resummation [17, 18]. To distinguish the scaling functions of $z$ introduced in Eq. (11), we use the bar notation for the scaling form of $\theta$. The normalization constants $\phi_{0}$ and $h_{0}$ are fixed by the conditions $1=\phi(h=+0,r=-1)=\phi(h=1,r=0)$, which reduces to
@@
We numerically solve Eq. (52) over a specific closed range of $Y$, yielding $\theta(h,\,r)$ and $R(h,\,r)$ denoted by $Z(Y)$. Then, we substitute $Z(Y)$ into (53) to obtain $\phi(Y)$. In Figs. 9(a)–9(c), we present the contours of $R$, $\theta$, and $\phi$, respectively. As observed from (a) and (b), $R$ ($R\geq 0$) represents the distance from the critical point at the origin, while $\theta$ ($-\theta_{\rm max}\leq\theta\leq\theta_{\rm max})$ measures the distance along the constant $R$ contour from the edges of the first-order boundary ($h=0,\ r<0$). Here, $\theta=\pm\theta_{\rm max}$ corresponds to $h=\pm 0$. With this parametrization, we can describe the criticality as a function of the $R$ variable, and the thermodynamic quantities will exhibit smooth behavior in the $\theta$ variable.
@@
We will determine the scaling function $\bar{p}(\theta)$ below to match with the relation between $\phi$ and $h$, Eq. (54). With Eq. (56) as a starting point, we can calculate the equations for the magnetization $\phi$ and the Ising energy $\varepsilon$ as (5):
@@
To derive the right-hand sides of Eq. (57), we use the derivative formula for an arbitrary function, $f\equiv R^{a_{f}}\bar{f}(\theta)$, with the scaling dimension $a_{f}$:
@@
which determines the form of $\bar{p}(\theta)$. Given that this is a first-order nonhomogeneous equation, it includes both special and general solutions. The special solution is relevant to our current analysis concerning the critical exponents beyond the mean-field level: $\bar{p}(\theta)=\sum_{n=0}^{3}C_{n}(1-\theta^{2})^{n}$, where $C_{n}$ are functions of the critical exponents and the parameters in Eq. (54) (for those explicit forms, refer to Ref. [3]). The constant of integration associated with the general solution $\propto(\theta^{2}-1)^{2-\alpha}$ is determined to be zero so that any singularity does not arise in the scaling part of the pressure, $\bar{p}(\theta)$.
@@
We highlight the qualitative distinction in behaviors for $\phi$ and $\varepsilon$, in Figs. 9(c) and 9(d), respectively. The former exhibits a monotonic increase as $\theta$ rises while keeping $R$ constant [also refer to Figs. 9(a) and 9(b) for tracking $R$ and $\theta$]. In contrast, the convex nature of $\varepsilon(\theta)$ results in a nonmonotonic ridge structure of $\varepsilon$. The distinctive ridgeline and the valley lines are illustrated in the magenta and cyan dashed lines according to the mathematical definition given by Eq. (104). Specifically, the ridgelines form a straight line with $\theta=0$ parametrized by an arbitrary $R>0$, due to symemtry of $\varepsilon$ under $h\rightarrow-h$ [^14] This is in contrast to $\hat{s}$ as we discuss at the end of Sec. IV.
@@
One can check that the the Maxwell relation (50) holds for the expressions (68b) and (68c), using Eq. (57).
@@
    We derived analytic expressions for $\phi(\theta,R)$, Eqs. (53) and (54a), and $\varepsilon(\theta,R)$, (66), respectively.
@@
    We numerically solved Eq. (52) with (54b), within a specified range of $h$ and $r$, and obtain $\theta(h,r)$ and $R(h,r)$.
@@
Here, the solution of the inverse function $\bar{h}_{2}^{-1}(z)$ restricted to (73) depends on the sign of $r$. If $r\leq 0$, then $\bar{h}_{2,-}^{-1}$ chose the solution of $\bar{h}_{2}^{-1}(z)$ satisfies $1\leq|\theta|<\theta_{\rm max}$; if $r>0$, then $\bar{h}_{2,+}^{-1}$ chose $10<|\theta|<$ 1. Using (74) on the $r\leq 0$ branch, we derive
@@
We can apply the same method to the Ising energy density $\varepsilon$. Eliminating $R$ from the scaling form in terms of $(\theta,R)$ variable, as given by (66), we have